OpenStudy (anonymous):

What is the perimeter of a parallelogram with sides of 2 and 3?

OpenStudy (anonymous):

2+2+3+3=

OpenStudy (anonymous):

Use the formula P= 2 ( L + W) P = 2 ( 2 + 3) P= 10